OpenStudy (anonymous):

sin 2x - 3 = -7 sin 2x find all solutions

OpenStudy (anonymous):

sin2x-3=-7sinx2 8sin2x=3 sin2x=3/8 Use sin inverse 2x=22 degrees, and 158 degrees Divide all by 2 x=11 degrees and 79 degrees

OpenStudy (anonymous):

thaaank you!